In how many ways can you form a four-digit number greater than or equal to 6000 using the digits 0,3,5,6,7, and 8 if repetition of digits is NOT allowed? 



1
Expert's answer
2021-03-03T05:59:02-0500

The first digit must be 6, 7 or 8.

Number of permutations from 5 by 3: "P_5^3=\frac{5!}{(5-3)!}=60."

Number of ways: "N=60*3=180."